# Volume polynomials

### v1

@@ 1. Realization problems for projection volumes and homology classes
Which tuples of six nonnegative real numbers can arise in this way? This question is the simplest nontrivial instance of the various realization problems for volume polynomials (Sections 2 and  3). The following answer was given in [HHM+, Theorem 1.4].
@@
Can you think of a simpler convex body in $\mathbb{R}^{4}$ that has the same six projection areas? See [HHM+, Section 4] for a discussion of this particular case.
@@
interpreted over the triangular hyperfield $\mathbb{T}_{2}$ on the set of nonnegative real numbers [BHKL]. As noted in [HHM+, Proposition 3.1], the condition is also equivalent to the statement that the nonnegative symmetric matrix
@@
there is no convex body $A$ in $\mathbb{R}^{5}$ such that the projection $\pi_{12}(A)$ has volume $4$ while all the other projections $\pi_{ij}(A)$ have volume $1$. On the other hand, every $4\times 4$ principal submatrix of the same matrix has at most one positive eigenvalue, suggesting that there is a convex body $A$ in $\mathbb{R}^{5}$ such that the projection $\pi_{345}(A)$ has area $4$ while all the other projections $\pi_{ijk}(A)$ have area $1$. Indeed, there is
@@
Which vectors of nonnegative integers $(p_{12},p_{13},p_{14},p_{23},p_{24},p_{34})$ can arise in this way? We call such homology classes realizable. [^1] Such questions are algebraic analogues of the Steenrod problem in topology [Eil49, Problem 25], which asks whether every homology class in any simplicial complex $X$ is the image of the fundamental class of a closed oriented manifold by a map into the simplicial complex. While the answer to this question is not known, the following partial result was given in [HHM+, Theorem 1.6].
@@
which has at most 2 solutions. This contradicts the assumption that $p_{34}=3$. The proof of Theorem 1.9 in [HHM+] shows that, in fact, the homology class corresponding to $(2,2,2,2,2,6)$ is realizable by an irreducible surface in $(\mathbb{P}^{1})^{4}$.
Modulo this subtlety involving integral coefficients, one can formulate statements in algebraic geometry parallel to Conjectures 1.6 and 1.7 in convex geometry, the first of which is a special case of [HHM+, Theorem 1.8].
@@ 2. Volume polynomials in convex geometry
In [Min03], Minkowski made a foundational observation that has since become a cornerstone of convex geometry:
@@
$f_{C}:\mathbb{R}^{n}_{\geq 0}\longrightarrow\mathbb{R}_{\geq 0},\qquad(x_{1},\ldots,x_{n})\longmapsto\operatorname{vol}(x_{1}C_{1}+\dots+x_{n}C_{n})$
@@
$\textrm{MV}(C_{i_{1}},\ldots,C_{i_{d}})\coloneqq\frac{1}{d!}\frac{\partial}{\partial x_{i_{1}}}\cdots\frac{\partial}{\partial x_{i_{d}}}f_{C}(x_{1},\ldots,x_{n}).$
@@
$\binom{d}{k}\,\textrm{MV}(I_{1},\ldots,I_{k},\underbrace{A,\dots,A}_{d-k})=\operatorname{vol}(\pi_{1\cdots k}A),$
where $\pi_{1\cdots k}$ is the projection onto the coordinate subspace orthogonal to $\mathbf{e}_{1},\ldots,\mathbf{e}_{k}$. For a comprehensive introduction to mixed volumes, see [Sch14, Chapter 5].
@@
The first condition is a formal consequence of the observation that the Minkowski sum of convex bodies is a convex body. The combination of the two properties leads to the conclusion that any volume polynomial must be a Lorentzian polynomial in the sense of [BH20]. Here we give an equivalent definition, following [BL, Section 2]. We set
@@
According to [BH20, Example 2.26], a bivariate polynomial with nonnegative coefficients
@@
In [She60], Shephard showed that any such polynomial is the volume polynomial of two convex bodies in $\mathbb{R}^{d}$. This characterizes volume polynomials of convex bodies in two variables:
@@
When every $p_{i}$ is rational, Shephard’s construction gives two rational convex polytopes. This is used in [Huh12, Theorem 21] to characterize realizable homology classes in $\mathbb{P}^{d}\times\mathbb{P}^{d}$ up to a multiple: Some nonnegative rational multiple of the class
@@
is the class of an irreducible subvariety if and only if $p_{0},\ldots,p_{d}$ is a log-concave sequence of nonnegative rational numbers with no internal zeros. [^2] As observed in [Huh, Section 5], there is no irreducible subvariety of $\mathbb{P}^{5}\times\mathbb{P}^{5}$ whose homology class corresponds to the log-concave sequence $(1,2,3,4,2,1)$.
@@
A quadratic form is Lorentzian if and only if all its coefficients are nonnegative and its Hessian has at most one positive eigenvalue [BH20, Section 2]. In [Hei38], Heine showed that, when there are at most three variables, any such quadratic form is the volume polynomial of three convex bodies in $\mathbb{R}^{2}$. This characterizes quadratic volume polynomials of convex bodies in three variables:
@@
The analogous statement fails when $n=4$. For example, as observed in [She60, Theorem 5], there are no convex bodies $C_{1},C_{2},C_{3},C_{4}$ in $\mathbb{R}^{2}$ satisfying
@@
even though the right-hand side is a Lorentzian polynomial. In fact, using the compactness theorem of Shephard for the affine equivalence classes of convex bodies [She60, Theorem 1], one can show that the displayed elementary symmetric polynomial is not even the limit of volume polynomials of convex bodies in the plane. This contrasts with the fact that there is an irreducible surface in $(\mathbb{P}^{1})^{4}$ with class $(1,1,1,1,1,1)$ in the Chow group. For example, one may take the closure of a general two-dimensional linear subspace of an affine chart of $(\mathbb{P}^{1})^{4}$.
@@
The main result of [BH20] provides a finite description of the set of Lorentzian polynomials that generalizes Example 2.2. The central notion is that of a generalized permutohedron. Let $E$ be a finite set with $n$ elements, and let $\{e_{i}\}_{i\in E}$ be the standard basis of $\mathbb{R}^{E}$.
@@
Generalized permutohedra are precisely the polytopes obtained from the standard permutohedron by moving the vertices so that all the edge directions are preserved [Pos09].
@@
The notion of $\mathrm{M}$-convex sets originates in discrete convex analysis [Mur03]. In [Mur03, Chapter 4], one can find several other equivalent characterizations of $\mathrm{M}$-convex sets. For example, a subset $J\subseteq\mathbb{Z}^{E}_{\geq 0}$ is $\mathrm{M}$-convex exactly when it satisfies the symmetric basis exchange property:
@@
For background specific to matroids, see [Oxl11].
@@
We recall the standard bijection between polymatroid rank functions on $E$ and nonempty $\mathrm{M}$-convex subsets of $\mathbb{Z}^{E}_{\geq 0}$ from [Mur03, Chapter 4]. For $A\subseteq E$ and $\alpha\in\mathbb{Z}^{E}_{\geq 0}$, we set $\alpha_{A}\coloneqq\sum_{i\in A}\alpha_{i}$.
@@
The subset $J(\varphi)$ is $\mathrm{M}$-convex for any $\varphi:E\to W$. Such matroids are said to be linear over $k$, and the function $\varphi$ is called a linear realization over $k$. One typically requires without loss of generality that the image of $\varphi$ spans $W$. A graphic matroid is linearly realizable over every field [Oxl11, Section 5.1]. In general, a matroid may or may not have a linear realization over $k$:
Among the four matroids pictured above, where the bases are given by all triples of points not on a line, the first is linear over $k$ if and only if the characteristic of $k$ is $2$, the second is linear over $k$ if and only if the characteristic of $k$ is not $2$, the third is linear over $k$ if and only if the cardinality of $k$ is not $2$, 33, or $5$, and the fourth is not linear over any field [Oxl11, Appendix].
@@
The subset $J(\varphi)$ is $\mathrm{M}$-convex for any $\varphi:E\to\ell$. Such matroids are said to be algebraic over $k$, and the function $\varphi$ is called an algebraic realization over $k$. One typically requires without loss of generality that the image of $\varphi$ contains a transcendence basis of $\ell$ over $k$. A linear matroid over $k$ is algebraic over $k$ [Oxl11, Section 6.7]. In general, a matroid may or may not have an algebraic realization over $k$:
Among the four matroids pictured above, where the bases are given by all triples of points not on a line, the first is algebraic over $k$ if and only if the characteristic of $k$ is $2$, the second and the third are algebraic over any field, and the fourth is algebraic over $k$ if and only if $k$ has nonzero characteristic [Oxl11, Appendix].
@@
where $\mathbb{M}^{d}_{n}\subseteq\mathbb{H}^{d}_{n}$ is the set of polynomials with nonnegative coefficients whose supports are $\mathrm{M}$-convex. [^3] By definition, the support of a polynomial $f$ is the set of all monomials appearing in $f$ with nonzero coefficients. The following characterization in [BH20, Theorem 2.25] is central to the theory of Lorentzian polynomials.
@@
$\small\left(\begin{array}[]{cccccc}0&1&1&\cdots&1\\ 1&0&1&\cdots&1\\ 1&1&0&\cdots&1\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&0\\ \end{array}\right),$
@@ 3. Volume polynomials in projective geometry
The analogous volume polynomial in algebraic geometry is defined as follows: Let $D=(D_{1},\ldots,D_{n})$ be a collection of semiample divisors on a $d$-dimensional projective variety $Y$ over a field kk. [^4] Throughout this paper, a variety over $k$ is by definition a reduced and irreducible scheme of finite type over $k$. A Cartier divisor on a complete variety is semiample if some positive multiple moves in a basepoint-free linear system. For background and any undefined terms concerning divisors on varieties, we refer to [Laz04]. The volume polynomial of $D$ is
@@
Also, by [Huh12, Theorem 21], for any $d$ and any $k$, we have
@@
As observed in [BH20, Section 4.2], it follows that
@@
A standard construction in toric geometry shows that any volume polynomial of rational convex polytopes arises as the volume polynomial of semiample divisors on projective varieties [Ful93, Section 5.4]. Thus, we have
@@
Since any convex body is a limit of a sequence of rational convex polytopes [Sch14, Section 1.8], we have
@@
Since volume polynomials of positive semidefinite Hermitian matrices are stable [Wag11, Proposition 2.1], the bivariate Lorentzian polynomial $x_{1}^{3}+6x_{1}^{2}x_{2}+6x_{1}x_{2}^{2}+2x_{2}^{3}$ is not in the left-hand side. On the other hand, any Lorentzian bivariate form is a volume polynomial over $k$ for any $k$, so the inclusion is strict. For a detailed discussion of volume polynomials of positive semidefinite Hermitian matrices, see [HMWX].
@@
where the sum is over all the $28$ bases of the Fano matroid. According to [GHM+, Example 5.5], we have
@@
If a matroid $J$ of rank $d$ on $n$ elements is linear over $k$, then its basis generating polynomial $f_{J}$ is a homogeneous polynomial of degree $d$ in $n$ variables. The arrangement Schubert variety of any linear realization of $J$ over $k$ witnesses the fact that $f_{J}$ is a realizable volume polynomial over $k$ [BHM+, Section 1.3].
@@
where $K_{\lambda\mu}$ is the Kostka number counting semistandard Young tableaux of given shape $\lambda$ and weight $\mu$ [Ful97, Section 8.3]. The normalized Schur polynomial is the generating polynomial
@@
The proof of [HMMSD22, Theorem 3] shows that any normalized Schur polynomial is a realizable volume polynomial over $k$ for any $k$.
@@
It is known that $\mathbb{V}^{d}_{n}(\mathbb{Q},k)$, and hence $\mathbb{V}^{d}_{n}(\mathbb{R},k)$, only depends on the characteristic of $k$ [GHM+, Proposition 2.10]. It is not known whether $\mathbb{V}^{d}_{n}(\mathbb{R},k)$ depends on $k$ when $d\geq 3$ and $n\geq 3$.
@@
As observed in [Huh23, Example 14], we have the proper inclusion
@@
each of which has exactly one positive eigenvalue. Then, by Theorem 2.11, $f$ is a Lorentzian polynomial. The fact that $f$ is not the volume polynomial over $k$ follows from the reverse Khovanskii–Teissier inequality [LX17, Theorem 5.7]: For any nef divisors $D_{1},D_{2},D_{3}$ on a $d$-dimensional projective variety $Y$ and any $e\leq d$,
${d\choose e}\ (D_{2}^{e}\cdot D_{1}^{d-e})_{Y}\ (D_{1}^{e}\cdot D_{3}^{d-e})_{Y}\geq(D_{1}^{d})_{Y}\ (D_{2}^{e}\cdot D_{3}^{d-e})_{Y}.$
The complex analytic proof of the inequality in [LX17] relies on the Calabi–Yau theorem [Yau78]. The algebraic proof of the inequality in [JL23] using Okounkov bodies works over any algebraically closed field. As mentioned before,
@@
Since Lorentzian polynomials are strongly log-concave [BH20, Theorem 2.31], the Lorentzian cubic $f$ provides a counterexample to Gurvits’ conjecture that a strongly log-concave homogeneous polynomial in three variables with nonnegative coefficients is the volume polynomial of three convex bodies [Gur09, Conjecture 4.1].
In [HMWX], the authors introduce a new family of inequalities for volume polynomials that subsumes both the Khovanskii–Teissier and the reverse Khovanskii–Teissier inequalities as special cases.
@@
When applied to the spanning tree polynomials of graphs and the basis generating polynomials of matroids, Definition 3.9 recovers the corresponding notions of contraction, deletion, and minor in the context of graph theory and matroid theory [Oxl11, Chapter 3]. [^5] In [Oxl11, Section 1.2], a matroid is required to have at least one basis. This leads to minor differences in the definition of $f\setminus j$ and $f/j$ when $j$ is a loop (when $j$ is contained in no basis of the matroid) or a coloop (when $j$ is contained in every basis of the matroid). For a discussion of minors in the more general framework of polymatroids over tracts, see [BKH+, Section 2.2].
A special case of [GHM+, Corollary 3.3] implies that deletions and contractions of realizable volume polynomials over $k$ are realizable volume polynomials over $k$. For a systematic study of linear operators preserving the set of realizable volume polynomials over $k$, see Section 4.
@@
By [Gro90], the $(1,1)$-part of the cohomology of $Y$ satisfies the mixed Hodge–Riemann relations, and hence
@@
It follows that the support of an analytic volume polynomial is $\mathrm{M}$-convex, defining the class of analytic polymatroids [GHM+, Section 5]. On the other hand, the resolution of singularities for complex projective varieties implies that any realizable volume polynomial over $\mathbb{C}$ is a realizable analytic volume polynomial, and hence

### v2

@@ 1. Realization problems for projection volumes and homology classes
Which tuples of six nonnegative real numbers can arise in this way? This question is the simplest nontrivial instance of the various realization problems for volume polynomials (Sections 2 and  3). The following answer was given in [20, Theorem 1.4].
@@
Can you think of a simpler convex body in $\mathbb{R}^{4}$ that has the same six projection areas? See [20, Section 4] for a discussion of this particular case.
@@
interpreted over the triangular hyperfield $\mathbb{T}_{2}$ on the set of nonnegative real numbers [3]. As noted in [20, Proposition 3.1], the condition is also equivalent to the statement that the nonnegative symmetric matrix
@@
there is no convex body $A$ in $\mathbb{R}^{5}$ such that the projection $\pi_{12}(A)$ has volume $4$ while all the other projections $\pi_{ij}(A)$ have volume $1$. On the other hand, every $4\times 4$ principal submatrix of the same matrix has at most one positive eigenvalue, suggesting that there is a convex body $A$ in $\mathbb{R}^{5}$ such that the projection $\varphi_{12}(A)$ has area $4$ while all the other projections $\varphi_{ij}(A)$ have area $1$. Indeed, there is
@@
Which vectors of nonnegative integers $(p_{12},p_{13},p_{14},p_{23},p_{24},p_{34})$ can arise in this way? We call such homology classes realizable. [^1] Such questions are algebraic analogues of the Steenrod problem in topology [11, Problem 25], which asks whether every homology class in any simplicial complex $X$ is the image of the fundamental class of a closed oriented manifold by a map into the simplicial complex. While the answer to this question is not known, the following partial result was given in [20, Theorem 1.6].
@@
which has at most 2 solutions. This contradicts the assumption that $p_{34}=3$. The proof of Theorem 1.9 in [20] shows that, in fact, the homology class corresponding to $(2,2,2,2,2,6)$ is realizable by an irreducible surface in $(\mathbb{P}^{1})^{4}$.
Modulo this subtlety involving integral coefficients, one can formulate statements in algebraic geometry parallel to Conjectures 1.6 and 1.7 in convex geometry, the first of which is a special case of [20, Theorem 1.8].
@@ 2. Volume polynomials in convex geometry
In [30], Minkowski made a foundational observation that has since become a cornerstone of convex geometry:
@@
$f_{C}:\mathbb{R}^{n}_{\geq 0}\longrightarrow\mathbb{R}_{\geq 0},\qquad(x_{1},\ldots,x_{n})\longmapsto\frac{1}{d!}\operatorname{vol}(x_{1}C_{1}+\dots+x_{n}C_{n})$
@@
$\textrm{MV}(C_{i_{1}},\ldots,C_{i_{d}})\coloneqq\frac{\partial}{\partial x_{i_{1}}}\cdots\frac{\partial}{\partial x_{i_{d}}}f_{C}(x_{1},\ldots,x_{n}).$
@@
$\binom{d}{k}\,\textrm{MV}(I_{1},\ldots,I_{k},\underbrace{A,\dots,A}_{d-k})=\frac{1}{k!}\operatorname{vol}(\pi_{1\cdots k}A),$
where $\pi_{1\cdots k}$ is the projection onto the coordinate subspace orthogonal to $\mathbf{e}_{1},\ldots,\mathbf{e}_{k}$. For a comprehensive introduction to mixed volumes, see [35, Chapter 5].
@@
The first condition is a formal consequence of the observation that the Minkowski sum of convex bodies is a convex body. The combination of the two properties leads to the conclusion that any volume polynomial must be a Lorentzian polynomial in the sense of [9]. Here we give an equivalent definition, following [10, Section 2]. We set
@@
According to [9, Example 2.26], a bivariate polynomial with nonnegative coefficients
@@
In [36], Shephard showed that any such polynomial is the volume polynomial of two convex bodies in $\mathbb{R}^{d}$. This characterizes volume polynomials of convex bodies in two variables:
@@
When every $p_{i}$ is rational, Shephard’s construction gives two rational convex polytopes. This is used in [24, Theorem 21] to characterize realizable homology classes in $\mathbb{P}^{d}\times\mathbb{P}^{d}$ up to a multiple: Some nonnegative rational multiple of the class
@@
is the class of an irreducible subvariety if and only if $p_{0},\ldots,p_{d}$ is a log-concave sequence of nonnegative rational numbers with no internal zeros. [^2] As observed in [23, Section 5], there is no irreducible subvariety of $\mathbb{P}^{5}\times\mathbb{P}^{5}$ whose homology class corresponds to the log-concave sequence $(1,2,3,4,2,1)$.
@@
A quadratic form is Lorentzian if and only if all its coefficients are nonnegative and its Hessian has at most one positive eigenvalue [9, Section 2]. In [18], Heine showed that, when there are at most three variables, any such quadratic form is the volume polynomial of three convex bodies in $\mathbb{R}^{2}$. This characterizes quadratic volume polynomials of convex bodies in three variables:
@@
The analogous statement fails when $n=4$. For example, as observed in [36, Theorem 5], there are no convex bodies $C_{1},C_{2},C_{3},C_{4}$ in $\mathbb{R}^{2}$ satisfying
@@
even though the right-hand side is a Lorentzian polynomial. In fact, using the compactness theorem of Shephard for the affine equivalence classes of convex bodies [36, Theorem 1], one can show that the displayed elementary symmetric polynomial is not even the limit of volume polynomials of convex bodies in the plane. This contrasts with the fact that there is an irreducible surface in $(\mathbb{P}^{1})^{4}$ with class $(1,1,1,1,1,1)$ in the Chow group. For example, one may take the closure of a general two-dimensional linear subspace of an affine chart of $(\mathbb{P}^{1})^{4}$.
@@
The main result of [9] provides a finite description of the set of Lorentzian polynomials that generalizes Example 2.2. The central notion is that of a generalized permutohedron. Let $E$ be a finite set with $n$ elements, and let $\{e_{i}\}_{i\in E}$ be the standard basis of $\mathbb{R}^{E}$.
@@
Generalized permutohedra are precisely the polytopes obtained from the standard permutohedron by moving the vertices so that all the edge directions are preserved [33].
@@
The notion of $\mathrm{M}$-convex sets originates in discrete convex analysis [31]. In [31, Chapter 4], one can find several other equivalent characterizations of $\mathrm{M}$-convex sets. For example, a subset $J\subseteq\mathbb{Z}^{E}_{\geq 0}$ is $\mathrm{M}$-convex exactly when it satisfies the symmetric basis exchange property:
@@
For background specific to matroids, see [32].
@@
We recall the standard bijection between polymatroid rank functions on $E$ and nonempty $\mathrm{M}$-convex subsets of $\mathbb{Z}^{E}_{\geq 0}$ from [31, Chapter 4]. For $A\subseteq E$ and $\alpha\in\mathbb{Z}^{E}_{\geq 0}$, we set $\alpha_{A}\coloneqq\sum_{i\in A}\alpha_{i}$.
@@
The subset $J(\varphi)$ is $\mathrm{M}$-convex for any $\varphi:E\to W$. Such matroids are said to be linear over $k$, and the function $\varphi$ is called a linear realization over $k$. One typically requires without loss of generality that the image of $\varphi$ spans $W$. A graphic matroid is linearly realizable over every field [32, Section 5.1]. In general, a matroid may or may not have a linear realization over $k$:
Among the four matroids pictured above, where the bases are given by all triples of points not on a line, the first is linear over $k$ if and only if the characteristic of $k$ is $2$, the second is linear over $k$ if and only if the characteristic of $k$ is not $2$, the third is linear over $k$ if and only if the cardinality of $k$ is not $2$, 33, or $5$, and the fourth is not linear over any field [32, Appendix].
@@
The subset $J(\varphi)$ is $\mathrm{M}$-convex for any $\varphi:E\to\ell$. Such matroids are said to be algebraic over $k$, and the function $\varphi$ is called an algebraic realization over $k$. One typically requires without loss of generality that the image of $\varphi$ contains a transcendence basis of $\ell$ over $k$. A linear matroid over $k$ is algebraic over $k$ [32, Section 6.7]. In general, a matroid may or may not have an algebraic realization over $k$:
Among the four matroids pictured above, where the bases are given by all triples of points not on a line, the first is algebraic over $k$ if and only if the characteristic of $k$ is $2$, the second and the third are algebraic over any field, and the fourth is algebraic over $k$ if and only if $k$ has nonzero characteristic [32, Appendix].
@@
where $\mathbb{M}^{d}_{n}\subseteq\mathbb{H}^{d}_{n}$ is the set of polynomials with nonnegative coefficients whose supports are $\mathrm{M}$-convex. [^3] By definition, the support of a polynomial $f$ is the set of all monomials appearing in $f$ with nonzero coefficients. The following characterization in [9, Theorem 2.25] is central to the theory of Lorentzian polynomials.
@@
$\left(\begin{array}[]{cccccc}0&1&1&\cdots&1\\ 1&0&1&\cdots&1\\ 1&1&0&\cdots&1\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&0\\ \end{array}\right),$
@@ 3. Volume polynomials in projective geometry
The analogous volume polynomial in algebraic geometry is defined as follows: Let $D=(D_{1},\ldots,D_{n})$ be a collection of semiample divisors on a $d$-dimensional projective variety $Y$ over a field kk. [^4] Throughout this paper, a variety over $k$ is by definition a reduced and irreducible scheme of finite type over $k$. A Cartier divisor on a complete variety is semiample if some positive multiple moves in a basepoint-free linear system. For background and any undefined terms concerning divisors on varieties, we refer to [27]. The volume polynomial of $D$ is
@@
Also, by [24, Theorem 21], for any $d$ and any $k$, we have
@@
As observed in [9, Section 4.2], it follows that
@@
A standard construction in toric geometry shows that any volume polynomial of rational convex polytopes arises as the volume polynomial of semiample divisors on projective varieties [12, Section 5.4]. Thus, we have
@@
Since any convex body is a limit of a sequence of rational convex polytopes [35, Section 1.8], we have
@@
Since volume polynomials of positive semidefinite Hermitian matrices are stable [37, Proposition 2.1], the bivariate Lorentzian polynomial $x_{1}^{3}+6x_{1}^{2}x_{2}+6x_{1}x_{2}^{2}+2x_{2}^{3}$ is not in the left-hand side. On the other hand, any Lorentzian bivariate form is a volume polynomial over $k$ for any $k$, so the inclusion is strict. For a detailed discussion of volume polynomials of positive semidefinite Hermitian matrices, see [22].
@@
where the sum is over all the $28$ bases of the Fano matroid. According to [16, Example 5.5], we have
@@
If a matroid $J$ of rank $d$ on $n$ elements is linear over $k$, then its basis generating polynomial $f_{J}$ is a homogeneous polynomial of degree $d$ in $n$ variables. The arrangement Schubert variety of any linear realization of $J$ over $k$ witnesses the fact that $f_{J}$ is a realizable volume polynomial over $k$ [8, Section 1.3].
@@
where $K_{\lambda\mu}$ is the Kostka number counting semistandard Young tableaux of given shape $\lambda$ and weight $\mu$ [13, Section 8.3]. The normalized Schur polynomial is the generating polynomial
@@
The proof of [21, Theorem 3] shows that any normalized Schur polynomial is a realizable volume polynomial over $k$ for any $k$.
@@
It is known that $\mathbb{V}^{d}_{n}(\mathbb{Q},k)$, and hence $\mathbb{V}^{d}_{n}(\mathbb{R},k)$, only depends on the characteristic of $k$ [16, Proposition 2.10]. It is not known whether $\mathbb{V}^{d}_{n}(\mathbb{R},k)$ depends on $k$ when $d\geq 3$ and $n\geq 3$.
@@
As observed in [25, Example 14], we have the proper inclusion
@@
each of which has exactly one positive eigenvalue. Then, by Theorem 2.11, $f$ is a Lorentzian polynomial. The fact that $f$ is not the volume polynomial over $k$ follows from the reverse Khovanskii–Teissier inequality [28, Theorem 5.7]: For any nef divisors $D_{1},D_{2},D_{3}$ on a $d$-dimensional projective variety $Y$ and any $e\leq d$,
$\binom{d}{e}\ (D_{2}^{e}\cdot D_{1}^{d-e})_{Y}\ (D_{1}^{e}\cdot D_{3}^{d-e})_{Y}\geq(D_{1}^{d})_{Y}\ (D_{2}^{e}\cdot D_{3}^{d-e})_{Y}.$
The complex analytic proof of the inequality in [28] relies on the Calabi–Yau theorem [39]. The algebraic proof of the inequality in [26] using Okounkov bodies works over any algebraically closed field. As mentioned before,
@@
Since Lorentzian polynomials are strongly log-concave [9, Theorem 2.31], the Lorentzian cubic $f$ provides a counterexample to Gurvits’ conjecture that a strongly log-concave homogeneous polynomial in three variables with nonnegative coefficients is the volume polynomial of three convex bodies [17, Conjecture 4.1].
In [22], the authors introduce a new family of inequalities for volume polynomials that subsumes both the Khovanskii–Teissier and the reverse Khovanskii–Teissier inequalities as special cases.
@@
When applied to the spanning tree polynomials of graphs and the basis generating polynomials of matroids, Definition 3.9 recovers the corresponding notions of contraction, deletion, and minor in the context of graph theory and matroid theory [32, Chapter 3]. [^5] In [32, Section 1.2], a matroid is required to have at least one basis. This leads to minor differences in the definition of $f\setminus j$ and $f/j$ when $j$ is a loop (when $j$ is contained in no basis of the matroid) or a coloop (when $j$ is contained in every basis of the matroid). For a discussion of minors in the more general framework of polymatroids over tracts, see [2, Section 2.2].
A special case of [16, Corollary 3.3] implies that deletions and contractions of realizable volume polynomials over $k$ are realizable volume polynomials over $k$. For a systematic study of linear operators preserving the set of realizable volume polynomials over $k$, see Section 4.
@@
By [15], the $(1,1)$-part of the cohomology of $Y$ satisfies the mixed Hodge–Riemann relations, and hence
@@
It follows that the support of an analytic volume polynomial is $\mathrm{M}$-convex, defining the class of analytic polymatroids [16, Section 5]. On the other hand, the resolution of singularities for complex projective varieties implies that any realizable volume polynomial over $\mathbb{C}$ is a realizable analytic volume polynomial, and hence